OpenStudy (anonymous):

i am unsure on how to find the x and y-intercepts of y=x^2-10x+21

OpenStudy (anonymous):

To find x intercepts, set y equal to zero (0=x^2-10x+12) and then factor the quadratic. You may need to use quadratic formula. To find the y intercepts, set the x equal to zero (y=0^2-10(0)+12) which just leaves y=12

OpenStudy (anonymous):

do i plug y=12 back into the equation to find the y-intercept? beacuse y=12 isn't an option on my assignment

OpenStudy (anonymous):

oops, that's a typo. I meant y=21. I transposed the digits

OpenStudy (anonymous):

and for the x-intercpet after i do the quadratic equation how do i find the other set of the coordinates?

OpenStudy (anonymous):

You'll probably end up with two answers from the quadratic formula. those should be your x-intercepts

OpenStudy (anonymous):

i did the formula and i got 

OpenStudy (kainui):

Take a second to consider what you're really looking at when you solve for the x and y intercepts. So where does the graph intersect the y-axis? When x=0! So just plug in 0 for x to solve for you y-intercept! The x-intercept? Similar thing! When y=0. Unfortunately you can't factor it easily but you can see from the graph that it will touch it in two different places, right? So using the quadratic formula you can find out where the graph touches those two points since the square root of any number can be either positive or negative. Don't forget that the square root of 4 isn't just 2 but you could also multiply -2 by itself to get 4!